OpenStudy (questionaskerduh1243):

Help Anyone? Will Medal And Fan! Look at the following numbers: −2, −1, 0, 2 Which pair of numbers has a sum of 0? −1, 2 −2, 0 −2, 2 −1, 0

OpenStudy (mathstudent55):

Opposites add to zero. Opposites are also called additive inverses.

OpenStudy (mathstudent55):

Here are some sums of additive inverses. 0 + 0 = 0 1 + (-1) = 0 10 + (-10) = 0 4.5 + (-4.5) = 0 Do you see any numbers among your given numbers that are additive inverses?